# Towards Exploring Continual Fine-Tuning for Enhancing Language Ability in Large Language Models

### v1

@@ Appendix C Mitigating Strategies
Here, we first visualize the impact of our mitigating strategies on the variance in model representations. Next, we provide ablate our findings for the Instruct-mOpenOrca phase-wise datasets.
Figure C4: Visualizing Variance in Model Representations for Mistral-7B Mitigating Strategies: We see a decrease in the variance of model representations for models trained using our mitigation strategies compared to vanilla Phase 2 models (see Figure 2).
Visualizing Variance in Model Representations.
In Figure C4, we repeat the same experiment as in § 4.5 to quantify the representation change in the fine-tuned models using our mitigating strategies. The trend seen is expected from §4.5: we see a decrease in the variation in the model activations, compared to the baseline model trained on Instruct and MultiAlpaca.
For the mitigating strategies that are curated to curb representational change, i.e., LF_H2, GR_5, and GR_10, we see that the corresponding curves have lesser change than the baseline Phase 2 model, Mistral-7B Instruct MultiAlpaca. That is, there is less representational change for LF_H2, GR_5, and GR_10 compared to Mistral-7B Instruct MultiAlpaca.
Our generative replay techniques are the closest in the representational change to Mistral-7B Instruct. This ‘closeness’ also improves its task and language ability performance compared to the vanilla Phase 2 model, Mistral-7B Instruct MultiAlpaca(refer to Table 1 and Table 2).
Additional Ablations.
@@
We also present the impact of our mitigating strategies for the Instruct-mOpenOrca phase-wise datasets on Mistral-7B. Table C6 presents these results.
@@
We see that LF_H2 achieves moderate success, especially in maintaining the language ability for MLQA (0.258) and XQUAD (0.527). However, task ability shows some decline (e.g., IFEval (0.401) and Alpaca Eval (0.048)), compared to the baseline. Furthermore, GR_5 results in lower task ability (IFEval = 0.281), while GR_10 performs slightly better in task ability (e.g., MMLU = 0.483, HellaSwag = 0.494).
Among the baselines, ER_10 performs similarly to the generative replay strategies, with modest improvements in task ability (e.g., IFEval = 0.367, MMLU = 0.479), but still struggles in language ability. Perhaps LoRA shows the best overall performance among the strategies for maintaining task ability (e.g., IFEval = 0.587, MMLU = 0.567, HellaSwag = 0.591) with reasonable retention of language ability (e.g., XQUAD = 0.354).
Note.
@@
CFT Setup Task Ability Language Ability Model Phase 1 Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Dataset Dataset Strategy Mistral-7B Instruct 0.55 0.35 0.575 0.641 0.529 0.246 0.012 0.351 0.203 mOpenOrca 0.426 0.06 0.507 0.509 0.376 0.155 0.040 0.323 0.173 LF_H2 0.401 0.048 0.518 0.487 0.364 0.258 0.060 0.527 0.282 GR_5 0.281 0.027 0.478 0.495 0.320 0.167 0.042 0.305 0.171 GR_10 0.305 0.013 0.483 0.494 0.324 0.150 0.038 0.238 0.142 LoRA 0.587 0.13 0.567 0.591 0.469 0.167 0.027 0.354 0.183 ER_10 0.367 0.025 0.479 0.493 0.341 0.157 0.042 0.305 0.168
Table C6: Task and Language ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10) and Layer Freezing (LF_H1 & LF_H2). We also use LoRA (hu2022lora) and ER_10 as two baseline strategies. Here, we perform Phase 2 with rank 64 and bf16 for LoRA. For ER_10, we use the English dataset used in GR_5 with original responses.
Appendix D Resources Used
We used 4 NVIDIA A100 GPU (80 GB) with a 96 core AMD CPU to run our inferences. One Finetuning Run with MultiAlpaca took 4 hours while for mOpenOrca it took 12 hours.
the list of model and the URL with checkpoints available and licenses are listed below:
  • LLaMA-3-8B : meta-llama/Meta-Llama-3-8B License: llama3
  • Mistral-7B : https://huggingface.co/mistralai/Mistral-7B-v0.1 License: Apache-2.0

### v3

@@ Appendix D Mitigating Strategies
Here, we provide additional details on Spectrum Hartford et al. (2024). We then visualize the impact of our mitigating strategies on the variance in model representations. Lastly, we ablate our findings for the Instruct-mOpenOrca phase-wise datasets.
D.1 Spectrum
Spectrum Hartford et al. (2024) is a layer-freezing technique that optimizes the fine-tuning of LLMs by selecting layers based on their signal-to-noise ratio (SNR). We use Spectrum as a heuristic for layer-freezing; that is, the layers identified as "important" by Spectrum are frozen during Phase 2 fine-tuning. A layer is important based on its signal-to-noise (SNR) ratio. In the following, we elaborate on how Spectrum computes SNR.
Marchenko-Pastur distribution.
The Marchenko-Pastur distribution Marchenko and Pastur (1967) is given by:
$\rho(\lambda)=\frac{1}{2\pi\sigma^{2}}\frac{\sqrt{(\lambda_{+}-\lambda)(\lambda-\lambda_{-})}}{\lambda},$
where
$\lambda_{\pm}=\sigma^{2}(1\pm\sqrt{Q})^{2},$
and $Q=\frac{N}{M}$, with $N$ and $M$ being the dimensions of a random matrix $W$, and $\sigma^{2}$ representing the variance of the entries in $W$.
SNR.
Let $W\in\mathbb{R}^{N\times M}$ be the weight matrix of a given layer. The empirical spectral density of $W$ is analyzed by comparing its eigenvalue distribution of $1/N\cdot W^{T}W$ against the theoretical Marchenko-Pastur distribution. Deviations from this distribution indicate the presence of significant signal components. We get,
$\lambda_{\pm}=\sigma^{2}\left(1\pm\sqrt{\frac{M}{N}}\right)^{2},$
where $\lambda_{\pm}$ are the largest and smallest eigenvalues and $\sigma$ the standard deviation. This implies the bounds of singular values of $W$ as:
$\epsilon_{\pm}=\frac{1}{\sqrt{N}}\sigma\left(1\pm\sqrt{\frac{M}{N}}\right)$ (1)
By evaluating how the singular values of $W$ distribute relative to $\epsilon_{\pm}$, Spectrum assesses the SNR of each layer, as defined next.
Ratio
Hartford et al. (2024). Specifically, the SNR value of a weight matrix is,
$\text{SNR}=\frac{\sum_{k|\sigma_{k}>\epsilon}\sigma_{k}}{\sum_{k|\sigma_{n}<\epsilon}\sigma_{n}}$
Here, $\epsilon$ separates signal from noisy singular values. Layers with singular values significantly exceeding $\epsilon_{+}$ have a high SNR, indicating a substantial presence of informative signal components.
Measuring the Ratio
Hartford et al. (2024). Having defined all ingredients above, Spectrum now computes each layer’s SNRs. To do this, it first computes SVD Zhang and Xu (2009) of the the layer’s weight matrix, calculates the SNR and normalizes it by the highest singular value. Eq. 1 gives the noise threshold.
Now, Spectrum selects layers with higher SNRs, where the number of layers selected is a hyperparameter. Similar to Hartford et al. (2024), for our experiments, we select the top-50% of layers in each module.
D.2 LLaMA-3-8B Doesn’t Show Consistent Improvement with our Mitigation Strategies
From Table 5, while both GR and LF improve on the baseline LLaMA-3-8B-Instruct MultiAlpaca, the gains in task and multilingual ability are not comparable to LLaMA-3-8B-Instruct.
(a) Mistral-7B
(b) LLaMA-3-8B
Figure D4: Demonstrating extent of cross-lingual transfer in Mistral-7B and LLaMA-3-8B on a parallel dataset prepared by subsampling Flores Costa-jussà et al. (2022). We find that the English activation cluster for LLaMA-3-8B is separated from the multilingual cluster, compared to Mistral-7B.
To understand this further, for GR, we investigate the cross-linguality difference between LLaMA-3-8B and Mistral-7B. Like Figure 1, we plot t-SNEs of the mean model activations for the Mistral-7B and LLaMA-3-8B base models on two parallel datasets, English and Multilingual. We create the parallel datasets by subsampling data from Flores Costa-jussà et al. (2022). In Figure D4, we see that the English activation cluster for LLaMA-3-8B is separated out from multilingual cluster, compared to Mistral-7B. This suggests that GR may not be as effective when the model has less cross lingual ability. While for LF, we acknowledge that our method to identify the layers to freeze may not be the best and better methods to identify which layers to freeze can be a direction for future work.
With regards to LLaMA-3-8B-Instruct MultiAlpaca MA results in Table 2, we believe that this is due to lack of cross-linguality in LLaMA-3-8B-Instruct and less data in MultiAlpaca which fails to cause sufficient representation drift to improve the model’s performance.
D.3 Additional Ablations
@@
We also present the impact of our mitigating strategies for the Instruct-mOpenOrca phase-wise datasets on Mistral-7B. Table D9 presents these results.
@@
We see that LF_H2 achieves moderate success, especially in maintaining the language ability for MLQA (0.258) and XQUAD (0.527). However, task ability shows some decline (e.g., IFEval (0.401) and Alpaca Eval (0.048)), compared to the baseline. Furthermore, GR_5 results in lower task ability (IFEval = 0.281), while GR_10 performs slightly better in task ability (e.g., MMLU = 0.483, HellaSwag = 0.494). Among the baselines, ER_10 performs similarly to the generative replay strategies, with modest improvements in task ability (e.g., IFEval = 0.367, MMLU = 0.479), but still struggles in language ability. Perhaps LoRA shows the best overall performance among the strategies for maintaining task ability (e.g., IFEval = 0.587, MMLU = 0.567, HellaSwag = 0.591) with reasonable retention of language ability (e.g., XQUAD = 0.354).
@@
CFT Setup Task Ability Language Ability Overall Model Phase 2 Mitigating IFEval Alpaca Eval MMLU HellaSwag Avg MLQA XLSum XQUAD Avg Avg Dataset Strategy Mistral-7B mOpenOrca 0.426 0.060 0.507 0.509 0.376 0.155 0.040 0.323 0.173 0.275 LF_H2 0.401 0.048 0.518 0.487 0.364 0.258 0.060 0.527 0.282 0.323 Spectrum 0.442 0.158 0.508 0.616 0.431 0.387 0.086 0.201 0.225 0.328 GR_5 0.281 0.027 0.478 0.495 0.320 0.167 0.042 0.305 0.171 0.246 GR_10 0.305 0.013 0.483 0.494 0.324 0.150 0.038 0.238 0.142 0.233 ER_10 0.367 0.025 0.479 0.493 0.341 0.157 0.042 0.305 0.168 0.255 LoRA 0.587 0.130 0.567 0.591 0.469 0.167 0.027 0.354 0.183 0.326
Table D9: English and Multilingual Ability results for our mitigating strategies, Generative Replay (GR_5 & GR_10), English Replay (ER_10) and Layer Freezing (LF_H1, LF_H2 & Spectrum). We use LoRA Hu et al. (2022) as a baseline strategy. For ER_10, we use the English dataset used in GR with original responses. The Phase 1 dataset is Instruct for each row. The first row provides Mistral-7B numbers for Instruct-mOpenOrca (from Table B2).
D.4 Compute Analysis
The computational overhead of replay arises from an increase in the effective training set size . In ER_10, each epoch includes the original Phase 2 data along with an additional 10% replay buffer from Phase 1, resulting in a proportional increase in the number of training tokens and hence compute. Similarly, GR_5 augments each epoch with 5% generated replay data, leading to a smaller overhead.
In contrast, layer freezing (LF) significantly reduces computation by updating only a subset of model parameters. By freezing 50% of the layers, LF yields a corresponding reduction in training cost while still achieving a reasonable trade-off between English retention and multilingual adaptation.
Appendix E Resources Used
We conduct all experiments on 4 NVIDIA A100 GPUs (80GB each) with a 96-core AMD CPU. A single fine-tuning run on MultiAlpaca takes approximately 4 hours, while training on mOpenOrca requires around 12 hours. The models used in our experiments, along with their corresponding checkpoints and licenses, are listed below:
Appendix F Existing Mitigating Strategies
We now discuss why MAD-X Pfeiffer et al. (2020) and other PEFT-based methods Badola et al. (2023) are not well-suited to our setting.
  • Parameter-Efficient Fine-tuning for Robust Continual Multilingual Learning Badola et al. (2023): This work considers a fixed downstream task (text classification) that is incrementally extended to new languages, mitigating forgetting via task- and language-specific adapters. Their approach assumes continued access to prior task data and the ability to maintain multiple adapter checkpoints. In contrast, our setting involves end-to-end instruction fine-tuning over a heterogeneous mixture of tasks. We do not assume access to Phase 1 data during Phase 2, nor do we maintain multiple parameter sets due to the memory overhead of storing additional model weights (§3).
  • MAD-X Pfeiffer et al. (2020): MAD-X introduces bottleneck adapters within encoder-only or encoder–decoder Transformer architectures to enable cross-lingual transfer. However, recent work Zhao et al. (2024) shows that such adapters do not integrate cleanly with decoder-only LLMs (e.g., Mistral-7B, LLaMA-3-8B), which form the basis of our experiments. As our models lack an encoder stack, the routing and merging mechanisms central to MAD-X are not directly applicable.
  • Preserving Cross-Linguality via Continual Learning Liu et al. (2021): This method extends Gradient Episodic Memory by storing exemplars (or their gradients) from previous phases and replaying them during training. Our setup (§3) explicitly disallows (i) reusing Phase 1 data and (ii) maintaining Phase 1 representations or weights during Phase 2, as these incur significant memory overhead and may violate practical data-sharing constraints.